If Y equals 10X plus 2 and 2X plus Y equals 26 what is the value of X minus Y
Found 2 solutions by richwmiller, ikleyn:
Answer by richwmiller(17219)

Try writing it as a math problem.
This is your lucky day.
Ikleyn decided to respond

Answer by ikleyn(52778)

.

    Y = 10X + 2     (1)

    2X + Y = 26     (2)


From equation(1), you just have an expression for Y via X.

Substitute this expression into equation (2). You will get then


    2X + (10x + 2) = 26.


Simplify and solve for X


    12X = 26 - 2

    12X = 24

      X = 24/12 = 2.


You just found X:  X = 2.


To find Y, substitute this value X= 2 into equation (2).  You will get


    Y = 10*2 +2 = 22.


Now you have everything to find the difference X-Y.


I leave this last step to you.

------------

To the tutor @richwmuller :

    I agree with you: for many visitors it is the way to get a solution and/or explanation that they practically 

    will get NEVER and from NOWHERE.


    I do it usually 10 - 20 times per day, when I find the problem formulation correct (or close to be correct)

    and then I think that my participation will be useful.


    Have a nice day (!)